OpenStudy (anonymous):

A hill has a grade of 6%. To the nearest degree, what is the angel that this hill makes with a horizontal line? Please help explain.

OpenStudy (aravindg):

u mean gradient?

OpenStudy (saifoo.khan):

let x be the angle so  tanx = 6/100 x= tan^-1 (6/100) so x = 3.433..... so that's 3 degrees